# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ 5. AI and FPGA Synergies: towards Next-Gen Earth Observation Missions
5.1. Cutting-Edge AI-FPGA Synergies for RS Applications (2.1)
While the setup of each study differs—making any comparison challenging and generalization even harder—we would like to highlight and discuss some of the best solutions surveyed. [^15] Dataset accuracies were compared to the benchmarks available at https://paperswithcode.com/sota.
Balancing Accuracy and Computation
Deploying large ML models on FPGAs often requires balancing task accuracy and hardware performance. In this regard, Ni et al. (2023) achieve competitive Overall Accuracy (OA) on the NWPU-RESISC45 dataset (Cheng et al., 2017) with both ResNet-34 ($93\%$) and VGG16 ($92\%$) deployed on their Flexible accelerator design. The authors leverage mixed-precision datatypes and multiples PEs to reach high throughput while achieving accuracies competitive with the State Of The Art (SOTA) of $96\%$. On DOTAv1.0 (Xia et al., 2018), Zhang et al. (2021) implement a lightweight version of YOLOv2, achieving $67\%$ mAP. Although the accuracy remains significantly below SOTA ($82\%$), their compressed YOLOv2 model enables on-the-edge computations. As a general note, comparing SOTA-driven studies with HW-aware FPGA implementations is challenging without a cost/benefit analysis. Quantifying the computational cost of a $1\%$ accuracy gain would clarify the trade-off between task performance and hardware efficiency, yet, this is often missing in SOTA literature.
Compact and Fast ML Models
Encoding models directly on-chip avoids loading parameters from peripherals, granting significant speed-up, as demonstrated by the OSCAR-RT framework in Yang et al. (2022). By mapping all CNN layers onto on-chip FPGA resources and using mixed-width quantization and optimized pipeline scheduling, they consistently compress YOLOv2 with different lightweight backbones under $0.5$MB while maintaining accuracies above $90\%$ mAP. In particular, their MobileNetv1 (Howard et al., 2017) and MobileNetv2 (Sandler et al., 2018) designs run in $0.7$ms, the lowest latency of the surveyed DL models. Wei et al. (2019) achieve LeNet-5 inference in $2.29$ms via a custom QAT algorithm, though their ablation study hints at sub-optimal FPGA resource usage. Upadhyay et al. (2024) classify cloud coverages with their SICNet CNN in $1.8$ ms, the lowest latency of Vitis AI implementations. The lowest per-pixel latency in this survey is achieved by the 2D CNN implemented by He et al. (2023), reaching $97\mu s$ on a Zynq US+. However, such performance is facilitated by the abundance of resources of the Zynq US+ series, which comes with a power consumption of $8.4$W.
Low-Power Implementations for Resource-Constrained Systems
As  2.1 highlights, power efficiency is the main concern for hardware designs, especially onboard SmallSats (Kothari et al., 2020). Traditional ML methods and shallow networks offer low power consumption, such as MLP (Hammoud et al., 2022), RDBC (Ratnakumar and Nanda, 2021), Fuzzy ARTMAP (Yahiaoui et al., 2017), and SVM-based models (Gyaneshwar and Nidamanuri, 2022) consuming less than $1$W. When it comes to DL, the CNN HO-ShipNet of Ieracitano et al. (2024) consumes $1.90$W on a Zynq 7045, while the Ghost-YOLOS of Yang et al. (2023) uses $2.98$W. Suh et al. (2021) achieve particularly impressive results using their custom 8-bit UniPOT quantization scheme and making use of DSP double rate, reaching $2.4$, $2.6$, and $2.0$W for their 3 SSD sizes. Among the solutions implemented with Vitis AI, Huang et al. (2024) reports the lowest power consumption, $1.65$W, a noteworthy result for a ResNet-50-based system like their ETAUS.
Lightweight Models and Quantization with Compensation
The combination of lightweight backbones and fixed-point quantization is a recurring and effective strategy in the surveyed studies (Yang et al., 2023, 2022; Yu et al., 2024; Li et al., 2023, 2019; Neris et al., 2022; Nguyen et al., 2024; Zhao et al., 2023; Kim et al., 2024). These two optimization methods are broadly applicable in RS, except in extreme scenarios where traditional ML methods are preferred to DNNs. Lightweight backbone replacement or direct use of an efficient architecture, like ENet (Paszke et al., 2016), or a ”tiny” adaptation, for example, YOLOv4-tiny (Yu et al., 2024; Nguyen et al., 2024), are used in $\approx 20\%$ of the surveyed experiments. To compensate for accuracy drops, designers typically enhance the baseline architecture. For example, Nguyen et al. (2024) add a third YOLOv4 head to detect targets at different spatial scales and optimize training with data augmentation strategies to achieve higher baseline accuracies. Following such enhancements, techniques like pruning or quantization can further reduce the solution’s computational load.
Accuracy loss due to design decisions during FPGA porting remains a challenge in achieving state-of-the-art results. While real-time performance is frequently achieved, most models still require significant power ($42\%$ $>5$W). Lightweight architectures combined with moderate ($>4$-bit) quantization offer a promising trade-off between accuracy and computational efficiency. 2.1 TAKEAWAYS
@@
Onboard PhiSat-1: Pioneering Work in AI for EO
PhiSat-11616166666U CubeSat, launched 09/2020, https://www.esa.int/Applications/Observing_the_Earth/Ph-sat is a frequently cited mission (Giuffrida et al., 2022) which deployed the first CNN model performing onboard cloud coverage classification: CloudScout (Giuffrida et al., 2020). ML-based methods accelerate data processing pipelines and recently became excellent at several essential tasks (Szwarcman et al., 2025). This is particularly true in RS, where the exponentially growing amount of data has outpaced the capacity for manual expert labeling. Deploying DL models directly onboard relieves downlink constraints and reduces the need for on-ground pre-processing. This saves time, conserves resources, and simplifies dataflows. For example, CloudScout classifies the cloud coverage of freshly acquired images, discarding overly cloudy images that contain little valuable EO information. Beyond hardware constraints, like limited memory and power, using a CNN for such a critical task implies a responsibility that comes with additional restrictions. As such, CloudScout has to ensure a False Positive rate under $1\%$, to avoid discarding appropriate images. In a follow-up study, Rapuano et al. (2021) explore the possibility of using an FPGA instead of the original Myriad 2 VPU. They conclude that, although FPGA-based designs take longer to develop and consume more power, their superior customizability, lower latency, and radiation hardness make them suitable for long-term LEO missions. In a recent work, Cratere et al. (2025) provide a comprehensive overview of the projects building on the CloudScout initiative, comparing all related cloud detection studies.
@@
A second ESA mission is particularly relevant to onboard ML for RS: OPS-SAT1717173333U CubeSat, 12/2019 - 05/2024, https://www.esa.int/Enabling_Support/Operations/OPS-SAT. This flying laboratory featured an Altera Cyclone V SoC and was designed for experimental ML deployments. The C-FCN++ model presented by Bahl et al. (2019) was deployed onboard OPS-SAT. With $273$ parameters ($47$KB), C-FCN++ is the only architecture explored in the study able to fit on-chip the Altera Cyclone V SoC. It uses dilated convolutions to expand early receptive fields and improve performance without increasing parameter count. Ultimately, C-FCN++ could perform cloud segmentation on a full-scale RS image in $150$ms, less than required for the camera of OPS-SAT to acquire the next image, granting real-time performance. For further investigations of ML onboard OPS-SAT, see Kacker et al. (2022), who tested several algorithms to prepare for the mission Beaver-Cube-21818183333U CubeSat educational mission from MIT, expected launch 2025, https://www.nanosats.eu/sat/beavercube-2.
@@
Spacecraft are not the only edge RS platforms benefiting from ML; numerous studies explore ML-supported pipelines for UAVs. Wang and Qiu (2019) deploy a CNN on a Zynq-7000 to detect objects in images from the 2018 DAC System Design Contest, targeting real-time inference under low-power constraints. Suh et al. (2021) design three sizes of SSD models to detect drones achieving high energy efficiency on a Zynq US+. Addressing a practical use case, Yu et al. (2024) develop an Improved YOLOv4-tiny to detect abnormal railway track fasteners. Implemented with Vitis AI, their solution achieves $\approx 300$ FPS and $95.1\%$ mAP, demonstrating performance levels suitable for onboard UAV deployment. Focusing on UAV navigation, Fraczek et al. (2018) experiment with Decision Trees and Support Vector Machines to classify terrain and support automated landing of UAVs. Similarly, Wang et al. (2024) merge optical and MilliMeter-Wave (MMW) radar data to quickly detect UAV obstacles. Their lightweight CNN reaches real-time performance with $60$ FPS and low-power $3.3$W on a Zynq-7020. By far the most complete surveyed UAV-focused study, Huang et al. (2024) present the ETAUS system, a custom UAV based on the Pixhawk 4 drone to monitor Taiwan’s Air Quality Index (AQI). ETAUS is powered by a CNN achieving high accuracy for AQI classification and cryptographic modules implemented on the FPGA logic. ETAUS also uses a YOLOv4 pre-trained from Vitis AI Model Zoo to detect—and later blur—private and sensitive information, such as license plates [^20] This section of the algorithm was not detailed in the study and is therefore not reported in this survey.. The study focuses on maintaining high accuracy for a reliable solution and achieving sufficient FPS for real-time performance.
@@
Real-time processing remains the central concern across the surveyed studies, as keeping pace with the continuous stream of newly acquired images is essential for operational success. To meet this demand, designers minimize model computations and maximize execution parallelism. In some contexts, such as the PhiSat-1 payload and CloudScout (Giuffrida et al., 2020), strict memory footprint limitations introduce additional constraints. Low power consumption is also a recurring objective, though the diversity of mission requirements makes it difficult to extract consistent trends. For instance, Nguyen et al. (2024) could afford to onboard a YOLOv4-tiny 3L with a $26.4$W power budget on their UAV.
@@ 6.1. Survey Limitations
While Scopus [^21] https://www.scopus.com/ and Web of Science (WoS) [^22] https://www.webofscience.com/ are the two most comprehensive public databases (Chadegani et al., 2013), we limited our search to WoS to maintain focus and feasibility; a preliminary analysis showed its journal coverage better aligns with the scope of this review. Besides, as with all systematic surveys, the search results inherently depend on the chosen query terms. This can lead to omissions, such as works deploying on FPGAs but using broader description like ’resource-limited hardware’ in their abstracts. Similarly, although we aimed for comprehensive ML-related queries, some lesser-known models may have been overlooked.
While the heterogeneity in the surveyed literature contributes to a broad and extensive body of work, it complicates direct comparisons of experiments. Comparing FPGA implementations is particularly challenging, as different designs use distinct networks and devices. Furthermore, most CNN-based studies do not benchmark against prior FPGA-based CNNs implementations, limiting our ability to identify commonly compared architectures. Lastly, distinguishing between Flexible and Specific accelerators often relies on our interpretation of the described designs.
@@ 6.2. Research Gaps Analysis
2.1 shows that the surveyed applications do not cover the full spectrum of RS problems. While edge AI does not resonate well with unstressed applications that can be processed on the ground, like long-term climate monitoring, other onboard-relevant use cases remain unexplored, such as data compression and disaster response. Indeed, the NewSpace era enables real-time alerts for numerous emergency situations, such as wildfires. Similarly, local authorities greatly benefit from post-disaster damage assessments, e.g., after floods (Mateo-Garcia et al., 2021). Exploring lightweight implementations of SOTA methods on FPGA could significantly advance both research and industry adoption. Concurrently, DL-based data compression pipelines have gained relevance over the past decade, including for EO data (Gomes et al., 2025). With SmallSats imaging payloads generating data at $\approx 640$ Mbps (Upadhyay et al., 2024), efficient onboard compression methods present a valuable research opportunity.
@@
Although  2.1 highlights the vast diversity of encountered models, the surveyed literature overlooks several widespread model families. Indeed, all DL-based studies rely on CNN or GNN architectures, with no examples of transformers or recurrent networks. This gap may partly reflect the recent emergence of Vision Transformers (ViTs), which have not yet seen mature FPGA implementations. However, it may also highlight the relative implementation simplicity of graph flows and convolution pipelines on FPGA platforms. Nevertheless, ViTs have significantly advanced DL in RS research (Szwarcman et al., 2025) and, despite their complexity, their superior performance makes them a promising target for edge deployment (Saha and Xu, 2025). Similarly, no surveyed work explores RNNs, despite their natural fit for temporal tasks—recurrent in RS, such as monitoring change across time or along flight paths. Investigating the application of RNNs onboard flying platforms represents a valuable research direction. At the same time, the absence of support for transformer and recurrent architectures in FPGA toolchains (e.g., FINN, Vitis AI) poses a barrier to broader adoption. To unlock the potential of modern DL architectures in edge environments, automated frameworks must evolve to support these emerging layers, including their quantized and pruned variants.
Beyond FPGAs: Coarse-Grained Reconfigurable Arrays (CGRA)
While  2.1 reveals a wide diversity of FPGAs, especially high-end products suited for DL inference, we fail to find ML solutions on CGRAs. Unlike FPGAs, CGRAs harden the commonly used arithmetic units, improving energy efficiency and allowing higher unit density. This reduces configurability, but offers enhanced performance. As discussed in Section 4.2, CNNs are often implemented with Processing Elements (PEs) organized in a systolic array. The AMD Versal [^23] https://www.amd.com/en/products/adaptive-socs-and-fpgas/versal.html platform features a systolic array of small PEs (AI Engines) capable of basic vector operation. Notably, the Versal AI Edge family combines FPGA programmable logic with CGRA AI engines for on-the-edge DNNs inference. Given the availability of space-grade Versal devices, a detailed comparison between CGRA-based and FPGA-only solutions–particularly for highly quantized models—emerges as a concrete and promising research direction.
@@
2.1 reveals that some widely used compression methods in Computer Vision remain underexplored in the surveyed records. In particular, Knowledge Distillation (KD) appears in only one study (Yang et al., 2023). KD is a training strategy where a large ’teacher’ model with high learning capabilities transfers its learned representations to a smaller ’student’ model (Hinton et al., 2015). In the on-the-edge context, the student network can be tailored to the target FPGA, maximizing resource utilization and performance. KD can also guide weight pruning to improve compression while limiting accuracy loss (Aghli and Ribeiro, 2021). This synergy holds strong potential for deployment in constrained environments. Although a few studies adopt hardware-aware design methods, hardware insights could benefit many optimization techniques. Notably, no surveyed work explores mixed-precision quantization driven by layer-wise hardware feedback, as in Yao et al. (2021), who improve energy efficiency and latency by assigning optimal bitwidths per layer. Similarly, we see potential in using AutoML, particularly Neural Architecture Search (NAS), to identify DNN architectures tailored to hardware constraints. Approaches like EfficientNet (Tan and Le, 2020) or MobiletNetv3 (Howard et al., 2019) are designed using platform-aware NAS, but hardware-aware NAS (HW-NAS) is still a young field with significant room for growth (Benmeziane et al., 2021). As GPUs increasingly integrate dedicated CNN units, such techniques become even more relevant for FPGAs, which must exploit their configurability to remain competitive. FPGA-aware NAS offers a promising path to design performant models well-matched to FPGA deployment, a research direction supported by datasets like HW-NAS-Bench (Li et al., 2021) that estimate FPGA performance.

### v2

@@ 5. AI and FPGA Synergies: towards Next-Gen Earth Observation Missions
In each experiment, designers must select an ML model, an FPGA device, and an accelerator architecture. In this vast search space, finding the best AI and FPGA technology combinations constitutes a tedious process. While the diversity of study setups makes any comparison challenging and generalization even more difficult, this section discusses promising associations between AI and FPGA technology.
5.1. Limits and suitabilities of AI-supported RS applications on FPGAs (2.1)
To demonstrate the feasibility of different ML models and RS applications for an increasing number of resources, we discuss trends among solutions based on the size of the FPGA. We use the amount of DSPs available, displayed in Fig. 4, as a proxy for the size of the FPGA as other resources (LUTs, FFs, BRAM) scale in a similar fashion.
Refer to caption Four lineplots displaying performance metrics per FPGA sizes The figure shows 4 line plots displaying 4 performance metrics on y-axes against the total number of DSPs available. The first plot displays the maximum utilization of DSPs and BRAM in %, and the line fitted using linear regression shows no correlation with the FPGA size. The second subplot presents the reported power, and the fit line grows steadily with larger FPGA devices. The third and fourth plots display the computational throughput and the memory board sizes, whose lines also grow with respect to the FPGA size.
Figure 5. Maximum utilization of DSP ’D[%] and BRAM ’B[%], reported power ’P[W], computational throughput [GOP/s], and footprint MB for different board sizes. We plot only reported metrics and fit the black line using linear regression.
Tiny and Small Devices
Among the four experiments using sub-150 DSP FPGAS, which we categorize as tiny, two designs deploy equally tiny models to perform pixel segmentation (Gyaneshwar and Nidamanuri, 2022) and image classification (Bahl et al., 2019). For its part, Vitolo et al. (2024) implement a larger network ($7403$ parameters) running at low frequency. All these designs achieve marginal power consumption (<0.5<0.5W) and demonstrate the use case of tiny FPGA devices for traditional ML models and tiny CNNs for fast approximate classification. The next size of FPGAs are small FPGA devices, often featured in educational evaluation boards. Two FPGAs of this class are used: first, the XC7Z020 with 220 DSPs and DDR3, and second, the newer generation XCZU3EG with 360 DSPs and faster DDR4 memory. Both devices successfully deploy traditional ML or shallow NN designs, but also smaller CNN-based models are implemented to segment images (Sabogal and George, 2021; Kim et al., 2024), classify ships (Hashimoto et al., 2019) or clouds (Pitonak et al., 2022), or detect miscellaneous objects (Wang and Qiu, 2019; Yang et al., 2023; Suh et al., 2021; Li et al., 2020b; Wang et al., 2024). Deploying CNNs on small FPGAs requires the use of reduced models (Yang et al., 2023; Pitonak et al., 2022; Wang et al., 2024) or sacrifices in FPS (Kim et al., 2024; Yang et al., 2023; Wang and Qiu, 2019; Hashimoto et al., 2019). Interestingly, two studies (Sabogal and George, 2021; Suh et al., 2021) examine the trade-off between model size, throughput, and accuracy. Suh et al. (2021) achieve particularly impressive results using their custom 8-bit UniPOT quantization scheme and leveraging DSP double rate, and perform inference <2.6<2.6W for all three SSD sizes.
Mid-Range FPGAs
FPGAs with more than 700 DSPs have sufficient resources for most classification or detection tasks using medium-sized models. However, to achieve high throughput, designers must still exploit significant model optimizations, such as heavy quantization, which leads to accuracy drops. For example, Rapuano et al. (2021) deploy a $\sim\!14$MB custom CNN with layer-specific QAT on an XCZU7EV and achieve inference in $141.7$ ms for 3.43.4W. Object detection tasks also induce a drop in accuracy, for example Zhang et al. (2021) implement a lightweight version of YOLOv2, achieving $67\%$ mAP on DOTAv1.0 (Xia et al., 2018), remaining significantly below the State Of the Art (SOTA) ($82\%$). Similarly, Zhao et al. (2023) leverage PTQ, pruning, and depth-wise separable convolutions to detect objects in the DIOR (Li et al., 2020a) dataset at $\sim\!50$FPS and under 7.27.2W. Mid-range devices can still provide low-power implementations, such as the NN HO-ShipNet (Ieracitano et al., 2024), consuming 1.901.90W on a Zynq 7045, or the ETAUS system (Huang et al., 2024), consuming 1.651.65W on a KV260.
High-End FPGAS
Only a few FPGAs incorporate 3500 DSPs or more. Such available computing power enables deployment without significant alteration of the models and facilitates results competitive with the SOTA [^15] Comparing SOTA-driven studies with HW-aware FPGA implementations is challenging without a cost/benefit analysis. Quantifying the computational cost of a $1\%$ accuracy gain would clarify the trade-off between task performance and hardware efficiency, yet this is often missing in SOTA literature.. On the relatively old Virtex-7, Ni et al. (2023) deploy large models for classification and object detection tasks. They achieve competitive accuracy on the NWPU-RESISC45 dataset (Cheng et al., 2017) with both ResNet-34 ($93\%$) and VGG16 ($92\%$) deployed on their Flexible accelerator design compared to the SOTA ($96\%$). As the DDR3 memory of the Virtex-7 can create a bottleneck on memory bandwidth, Yang et al. (2022) stores model weights on-chip. Through pruning and mixed-precision QAT, they consistently compress YOLOv2 with different lightweight backbones to under 0.50.5MB while maintaining high accuracies. Their model optimizations, paired with optimized pipeline scheduling and the use of $\sim\!70\%$ of the available DSPs, lead to the lowest latencies of the surveyed DL models, achieving 0.70.7ms for their MobileNetv1 (Howard et al., 2017) and MobileNetv2 (Sandler et al., 2018) designs. Two studies experiment with Alveo cards (Zhang et al., 2023; Wickramasinghe et al., 2024) and demonstrate impressive throughput performance, including for the only transformer architecture of this survey. However, datacenter-grade boards, such as the Alveo family, have power consumptions ranging from $75$ to 225225W, making them unsuitable for edge operations.
Synergizing Model and FPGA
Several factors influence model selection, foremost the hardware constraints discussed above and in  2.1. Indeed, limited computational units (DSPs, LUTs) and on-chip storage (BRAMs) impose a trade-off between accuracy and computational throughput. However, as seen in Fig. 5, resource utilization varies significantly across all FPGA sizes. Optimizing utilization could make smaller FPGAs viable for a broader range of applications. Additionally, the characteristics of ML tasks and RS applications also restrict model selection. For example, object detection problems on the edge typically rely on single-shot models, such as YOLO or SSD, to save compute time. For segmentation, patch-based models (e.g., UNet) tend to yield better task accuracy (Asgari Taghanaki et al., 2021) than methods that process pixels one at a time. Such models, however, have a larger memory footprint and higher computational costs. Classification, on the other hand, can be tackled with diverse models, ranging from SVMs that fit on all FPGA devices to larger models like ResNet50. As a final factor, because automatic toolchains only support a limited set of operations, designers who desire to accelerate development time and reduce the required design expertise must favor standard layers in their architectures. Ultimately, optimal model selection depends on each project’s performance requirements, deployment constraints, and available development effort.
Larger FPGAs naturally support higher model complexity and throughput, reducing the need for aggressive quantization or pruning. While real-time performance is frequently achieved, most models still require significant power ($43\%$ >5>5W). High variance in resource utilization shows that efficient resource use, rather than raw device size, is a key determinant of achievable performance. 2.1 TAKEAWAYS
@@
Onboard PhiSat-1: Pioneering Work in AI for RS
PhiSat-116161666U CubeSat, launched 09/2020, https://www.esa.int/Applications/Observing_the_Earth/Ph-sat is a frequently cited mission (Giuffrida et al., 2022) which deployed the first CNN model performing onboard cloud coverage classification: CloudScout (Giuffrida et al., 2020). ML-based methods accelerate data processing pipelines and recently became excellent at several essential tasks (Szwarcman et al., 2025). This is particularly true in RS, where the exponentially growing amount of data has outpaced the capacity for manual expert labeling. Deploying DL models directly onboard relieves downlink constraints and reduces the need for on-ground pre-processing. This saves time, conserves resources, and simplifies dataflows. For example, CloudScout classifies the cloud coverage of freshly acquired images, discarding overly cloudy images that contain little valuable information. Beyond hardware constraints, like limited memory and power, using a CNN for such a critical task implies a responsibility that comes with additional restrictions. As such, CloudScout has to ensure a False Positive rate under $1\%$, to avoid discarding appropriate images. In a follow-up study, Rapuano et al. (2021) explore the possibility of using an FPGA instead of the original Myriad 2 VPU. They conclude that, although FPGA-based designs take longer to develop and consume more power, their superior customizability, lower latency, and radiation hardness make them suitable for long-term LEO missions. In a recent work, Cratere et al. (2025) provide a comprehensive overview of the projects building on the CloudScout initiative, comparing all related cloud detection studies.
@@
A second ESA mission is particularly relevant to onboard ML for RS: OPS-SAT17171733U CubeSat, 12/2019 - 05/2024, https://www.esa.int/Enabling_Support/Operations/OPS-SAT. This flying laboratory featured an Altera Cyclone V SoC and was designed for experimental ML deployments. The C-FCN++ model presented by Bahl et al. (2019) was deployed onboard OPS-SAT. With $273$ parameters (4747KB), C-FCN++ is the only architecture explored in the study able to fit on-chip the Altera Cyclone V SoC. It uses dilated convolutions to expand early receptive fields and improve performance without increasing parameter count. Ultimately, C-FCN++ could perform cloud segmentation on a full-scale RS image in 150150ms, less than required for the camera of OPS-SAT to acquire the next image, granting real-time performance. For further investigations of ML onboard OPS-SAT, see Kacker et al. (2022), who tested several algorithms to prepare for the mission Beaver-Cube-218181833U CubeSat educational mission from MIT, expected launch 2025, https://www.nanosats.eu/sat/beavercube-2.
@@
Spacecraft are not the only edge RS platforms benefiting from ML; numerous studies explore ML-supported pipelines for UAVs. Wang and Qiu (2019) deploy a CNN on a Zynq-7000 to detect objects in images from the 2018 DAC System Design Contest, targeting real-time inference under low-power constraints. Suh et al. (2021) design three sizes of SSD models to detect drones achieving high energy efficiency on a Zynq US+. Addressing a practical use case, Yu et al. (2024) develop an Improved YOLOv4-tiny to detect abnormal railway track fasteners. Implemented with Vitis AI, their solution achieves $\sim\!300$ FPS and $95.1\%$ mAP, demonstrating performance levels suitable for onboard UAV deployment. Focusing on UAV navigation, Fraczek et al. (2018) experiment with Decision Trees and Support Vector Machines to classify terrain and support automated landing of UAVs. Similarly, Wang et al. (2024) merge optical and MilliMeter-Wave (MMW) radar data to quickly detect UAV obstacles. Their lightweight CNN reaches real-time performance with $60$ FPS and low-power 3.33.3W on a Zynq-7020. By far the most complete surveyed UAV-focused study, Huang et al. (2024) present the ETAUS system, a custom UAV based on the Pixhawk 4 drone to monitor Taiwan’s Air Quality Index (AQI). ETAUS is powered by a CNN achieving high accuracy for AQI classification and cryptographic modules implemented on the FPGA logic. ETAUS also uses a YOLOv4 pre-trained from Vitis AI Model Zoo to detect—and later blur—private and sensitive information, such as license plates [^20] This section of the algorithm was not detailed in the study and is therefore not reported in this survey.. The study focuses on maintaining high accuracy for a reliable solution and achieving sufficient FPS for real-time performance.
@@
Real-time processing remains the central concern across the surveyed studies, as keeping pace with the continuous stream of newly acquired images is essential for operational success. To meet this demand, designers minimize model computations and maximize execution parallelism. In some contexts, such as the PhiSat-1 payload and CloudScout (Giuffrida et al., 2020), strict memory footprint limitations introduce additional constraints. Low power consumption is also a recurring objective, though the diversity of mission requirements makes it difficult to extract consistent trends. For instance, Nguyen et al. (2024) could afford to onboard a YOLOv4-tiny 3L with a 26.426.4W power budget on their UAV.
@@ 6.1. Survey Limitations
While Scopus [^21] https://www.scopus.com/ and Web of Science (WoS) [^22] https://www.webofscience.com/ are the two most comprehensive public databases (Chadegani et al., 2013), we limited our search to WoS to maintain focus and feasibility; a preliminary analysis showed its journal coverage better aligns with the scope of this review. Besides, as with all systematic surveys, the search results inherently depend on the chosen query terms. This can lead to omissions, such as works deploying on FPGAs but using broader description like ’resource-limited hardware’ in their abstracts. Similarly, although we aimed for comprehensive ML-related queries, some lesser-known models may have been overlooked. While the heterogeneity in the surveyed literature contributes to a broad and extensive body of work, it complicates direct comparisons of experiments. Comparing FPGA implementations is particularly challenging, as different designs use distinct networks and devices. Furthermore, most CNN-based studies do not benchmark against prior FPGA-based CNNs implementations, limiting our ability to identify commonly compared architectures. Lastly, distinguishing between Flexible and Specific accelerators often relies on our interpretation of the described designs.
@@ 6.2. Research Gaps Analysis
2.1 shows that the surveyed applications do not cover the full spectrum of RS problems. While edge AI does not resonate well with unstressed applications that can be processed on the ground, like long-term climate monitoring, other onboard-relevant use cases remain unexplored, such as data compression and disaster response. Indeed, the NewSpace era enables real-time alerts for numerous emergency situations, such as wildfires. Similarly, local authorities greatly benefit from post-disaster damage assessments, e.g., after floods (Mateo-Garcia et al., 2021). Exploring lightweight implementations of SOTA methods on FPGA could significantly advance both research and industry adoption. Concurrently, DL-based data compression pipelines have gained relevance over the past decade, including for RS data (Gomes et al., 2025). With SmallSats imaging payloads generating data at $\approx 640$ Mbps (Upadhyay et al., 2024), efficient onboard compression methods present a valuable research opportunity. In addition to these underexplored use cases, retrieval problems remain severely underrepresented, with only two studies focusing on feature estimation, despite these tasks being frequently essential for practical remote sensing outcomes.
@@
Although  2.1 highlights the vast diversity of encountered models, the surveyed literature overlooks several widespread model families. Indeed, almost all DL-based studies rely on CNN or GNN architectures, with a single example of transformer models (Wickramasinghe et al., 2024). This gap may partly reflect the recent emergence of Vision Transformers (ViTs), which have not yet seen mature FPGA implementations. However, it may also highlight the relative implementation simplicity of graph flows and convolution pipelines on FPGA platforms. Nevertheless, ViTs have significantly advanced DL in RS research (Szwarcman et al., 2025) and, despite their complexity, their superior performance makes them a promising target for edge deployment (Saha and Xu, 2025). Similarly, no surveyed work explores recurrent models, despite their natural fit for temporal tasks—common in RS, such as monitoring change across time or along flight paths. Investigating the application of RNNs onboard flying platforms represents a valuable research direction. At the same time, the absence of support for transformer and recurrent architectures in FPGA toolchains (e.g., FINN, Vitis AI) poses a barrier to broader adoption. To unlock the potential of modern DL architectures in edge environments, automated frameworks must evolve to support these emerging layers, including their quantized and pruned variants.
Beyond FPGAs: Coarse-Grained Reconfigurable Arrays (CGRAs)
While we discovered a wide range of FPGAs used for RS applications, we found limited use of CGRAs. Unlike FPGAs, CGRAs harden the commonly used arithmetic units, improving energy efficiency and allowing higher density. This reduces configurability, but offers enhanced performance and power efficiency. As discussed in Section 4.2, CNNs are often implemented with Processing Elements (PEs) organized in a systolic array. The AMD Versal [^23] https://www.amd.com/en/products/adaptive-socs-and-fpgas/versal.html platform features a systolic array of small PEs (AI Engines) capable of basic vector operation. Perryman et al. (2025) uses two Versal devices (AI Edge and AI Core) to implement semantic segmentation using distinct networks. Compared to Sabogal and George (2021), which uses the same U-Net model and dataset (ISPRS, 2022), it increases the $\frac{FPS}{W}$ by 2.2x and 4.1x for the Edge and Core devices, respectively, demonstrating the potential of CGRAs. Given the availability of space-grade Versal devices, a detailed comparison between CGRA-based and FPGA-only solutions—particularly for highly quantized models—emerges as a concrete and promising research direction.
@@
2.1 reveals that some widely used compression methods in Computer Vision remain underexplored in the surveyed records. In particular, Knowledge Distillation (KD) appears in only one study (Yang et al., 2023). KD is a training strategy where a large ’teacher’ model with high learning capabilities transfers its learned representations to a smaller ’student’ model (Hinton et al., 2015). In the on-the-edge context, the student network can be tailored to the target FPGA, maximizing resource utilization and performance. KD can also guide weight pruning to improve compression while limiting accuracy loss (Aghli and Ribeiro, 2021). This synergy holds strong potential for deployment in constrained environments. Although a few studies adopt hardware-aware design methods, hardware insights could benefit many optimization techniques. Notably, no surveyed work explores mixed-precision quantization driven by layer-wise hardware feedback, as in Yao et al. (2021), who improve energy efficiency and latency by assigning optimal bitwidths per layer. Similarly, we see potential in using AutoML, particularly Neural Architecture Search (NAS), to identify DNN architectures tailored to hardware constraints. Approaches like EfficientNet (Tan and Le, 2020) or MobiletNetv3 (Howard et al., 2019) are designed using platform-aware NAS, but hardware-aware NAS (HW-NAS) is still a young field with significant room for growth (Benmeziane et al., 2021). In this survey, only Hammoud et al. (2024) used an approach similar to HW-NAS to determine the best model for their application. As GPUs increasingly integrate dedicated CNN units, such techniques become even more relevant for FPGAs, which must exploit their configurability to remain competitive. FPGA-aware NAS offers a promising path to design performant models well-matched to FPGA deployment, a research direction supported by datasets like HW-NAS-Bench (Li et al., 2021) that estimate FPGA performance.